40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 5352     Entry time: Wed Sep 7 00:39:34 2011
Author: rana 
Type: Update 
Category: SUS 
Subject: ITMX adjustments 

(What we did)

* Moved SUS to edge of table for OSEM adjustment.

* Leveled the table in this temporary tower position.

* Rotated all OSEMs to give some seperation between magnets and LED/PD packages.

* Moved the upper OSEM bracket a little bit upward.

* All the OSEM holding set screws were short with flat heads; this is annoying since we would like to use them more like thumbscrews. Steve took the long set-screws out of the old ITMX cage and we swapped them. Need to order ~100 silver-plated socket head spare/replacements.

* Took pictures of OSEMs.

* Moved tower back to old position.

* Releveled the table (added one rectangular weight in the NW corner of the table).

* Find that ITMX OSEMs were a couple 100 micron out of position; we adjusted them in-situ in the final position of the tower, trying not to rotate them. All mean voltages now are within 100 mV of ideal half-light.

* Back/front EQ positions adjusted by the screw method. bottom/top stops adjusted earlier.

* OSEM cables tied down with copper wire.

* Increased the incident power up to 91 mW going into MC to temporarily make the POX beam more visible.

* The POX beam was checked. It was exiting from the chamber and going through about the center of the viewport.

ELOG V3.1.3-